A person spent rs 564 buying pens and pencils. If the cost of each pen is rs 7 and each pencil is rs 3 and if the total number of things bought was 108, how many of each did he buy?
Found 2 solutions by Cromlix, macston:
Answer by Cromlix(4381)

Hi there,
Pen cost rs7
Pencil cost rs3
Total number of things = 108
The person spent rs 564
Make pens = x
Make pencils = y
x + y = 108......(1)
7x + 3y = 564.....(2)
Multiply (1) by 7
7x + 7y = 756.....910
7x + 3y = 564
Subtract (2) from (1)
....4y = 192
....y = 48
Substitute y = 48
into Equation (1)
x + y = 108
x + 48 = 108
x = 108 = 48
x = 60.
Therefore:
48 pencils were bought
60 pens were bought.
Hope this helps:-)

Answer by macston(5194)

.
X=number of pens; Y=number of pencils
.
X+Y=108
Y=108-X Use this to substitute for Y below.
.
rs7X+rs3Y=rs 564 Substitute for Y from above.
rs7X+rs3(108-X)=rs 564
rs7X+rs324-rs3X=rs 564 subtract rs 324 from each side.
rs4X=rs 240 Divide each side by rs 4.
X=60 ANSWER 1: The person bought 60 pens.
.
Y=108-X
Y=108-60
Y=48 ANSWER 2: The person bought 48 pencils.
.
CHECK:
A person spent rs 564 buying pens and pencils.
rs7X+rs3Y=rs 564
rs7(60)+rs3(48)=rs 564
rs 420+rs 144=rs 564
rs 564=rs564